Estimate the mass of human head.
Use Fermi calculations.
1
Expert's answer
2021-10-10T15:56:04-0400

The mass is given as follows:


"m = \rho V"

where "\rho" is the density of a head, and "V" is its volume. Let's approximate the head as a perfect shere. Then its volume is approximately 4 times its radius cubed or 0.5 times diameter cubed (assuming "\pi\approx 3" ):


"V \approx 0.5d^3"

Let's estimate the diameter of the head as "d\approx 20cm = 0.2m". Also, since the head consists mostly of brain (which has almost the same density as water) and bones (which are a little more dense than water), let's take the average density to be "\rho \approx 1200 kg/m^3".

Finally, obtain:


"m\approx 0.5d^3\rho\approx 0.5\cdot (0.2m)^3\cdot 1200kg/m^3 \approx 4.8kg"

The actual masses varies from approximately 4kg to 6 kg (see, for example, https://www.sciencedirect.com/science/article/abs/pii/S0021929009001444), thus, the obtainded estimation is quite good.


Answer. 4.8 kg.